# ControlVAR: Exploring Controllable Visual Autoregressive Modeling

### v1

@@ 4.1 Evaluation Settings
We conduct all the experiments on the ImageNet [88] dataset. To incorporate pixel-level controls, we leverage state-of-the-art image understanding models to pseudo-label the images. Specifically, we label entity masks [89], canny [90], depth [91] and normal [92] for both training and validation sets. This takes 500 Tesla V100 for about 4 days. We will release the pseudo-labeled datasets to facilitate the community to further explore conditional image generation.
@@
We utilize Fréchet Inception Distance (FID) [93], Inception Score (IS) [94], Precision, and Recall as metrics for assessing the quality of image generation. However, for the image-to-control prediction where ground truth is unavailable, we rely on qualitative visualization to demonstrate the perceptual quality.
@@
We follow VAR [4] to use a GPT-2 [70] style transformer with adaptive normalization [95]. A transformer layer depth from 12 to 30 is explored. We leverage the pre-trained VAR tokenizer [4] to tokenize both image and control. We initialize the model with the weights from VAR [4] to shorten the training process. For each depth, we train the model for 30 epochs with an Adam optimizer. We follow the same learning rate and weight decay as VAR. During training, we sample each control type uniformly. To apply the classifier-free guidance, we replace class and control type conditions with empty tokens with 0.1 probability. We train the model with $\textrm{batchsize}=128$ for all the experiments. During inference, we utilize top-$k$ top-$p$ sampling with $k=900$ and $p=0.96$. We utilize $256\times 256$ image size for all experiments. For simplicity, we leverage $\gamma_{cls}=\gamma_{typ}=\gamma_{pix}$ for all the experiments.
@@ 4.2 Performance Analysis
Furthermore, we compare the image FID with pure image generation model VAR [4] in Tab 1. We notice that ControlVAR shows a slight performance degradation compared to VAR which can be due to the difficulty enrolled to incorporate additional controls. As the model size increases, we notice the performance gap shrinks, indicating joint modeling of image and control may require more network capacity compared to image-only modeling.
Figure 7: Visualization of conditional image impainting (given pixel-level control and partial image).
@@
We introduce two baseline methods - ControlNet [1] and T2I-Adapter [2] to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser [96] implementation (for a fair comparison, ImageNet-pretrained LDM [6] is used as the base model). We compare the image generation quality in terms of FID and speed in Fig. 8. We notice that ControlVAR achieves obvious superior FID compared to baselines. We evaluate the inference speed with $\mathrm{batchsize}=1$ on a single H100 GPU. We notice that ControlVAR inference is at least 5 times faster than the compared methods. We further explore the generation capability with different model sizes as shown in Fig. 8 (b), we notice that the model’s generation capability keeps improving as the model size increases. Similar to the joint image-control generation, we notice that canny-conditioned generation shows an inferior performance due to its complex pattern.
@@
We further compare ControlVAR with conditional AR models - VQ-GAN [3] and IQ-VAE [24]. We finetune ControlVAR on ADE 20K for 1 epoch and report the FID of the generated images in Tab 2. ControlVAR demonstrates superior performance compared to previous AR methods.
@@ 5 Conclusion
In spite of ControlVAR’s high performance on image generation with heterogeneous pixel-level controls, it does not support text prompts and therefore cannot be directly leveraged with natural language guidance. Developing text-guided capability can be achieved by replacing the class token with the language token, e.g., CLIP token [97], which is left as our future focus.
@@ Appendix A Dataset Creation
We conduct all the experiments on the ImageNet [88] dataset. To incorporate pixel-level controls, we leverage state-of-the-art image understanding models to pseudo-label the images. Specifically, we label entity masks [89], canny [90], depth [91] and normal [92] for both training and validation sets. This takes 500 Tesla V100 for about 4 days. We demonstrate the label number after filtering in Tab A. In addition, we also manually check the quality of the pseudo labels. We show a visualization of the generated datasets in Fig. A. We notice that image understanding models predict reasonable results on ImageNet images.
@@ Appendix B Discussion of the Teacher Forcing Guidance
Inspired by the classifier-free guidance [13] from diffusion models, we empirically find a similar form of guidance that can be used for autoregressive sample conditional images based on teacher forcing. In this section, we aim to analyze the spirit of classifier-free guidance (CFG) and analogy it to our teacher-forcing guidance (TFG).
@@ B.1 Classifier-free Guidance
where $\gamma_{C},\gamma_{c},\gamma_{c_{t}}$ are the guidance scales that are used to adjust the amplitude to apply the conditional guidance. $\emptyset$ denotes leveraging a special empty token to replace the original token to disable the additional conditional information [13].
@@ B.2 Teacher Forcing Guidance
Classifier-free guidance has been proven to be effective in AR models [7, 4] which take the same form as diffusion models as
@@
where $p(I,C|c,c_{t})$ and $p(I,C|c,c_{t})$ can be found from the output of ControlVAR. We follow previous works [4, 3] to ignore the constant probabilities $p(c)$. By rewriting all terms with Baysian rule, we have
@@ C.1 Details of Evaluation Metrics
Fréchet Inception Distance (FID) [93]. FID measures the distance between real and generated images in the feature space of an ImageNet-1K pre-trained classifier [98], indicating the similarity and fidelity of the generated images to real images.
Inception Score (IS) [94]. IS also measures the fidelity and diversity of generated images. It consists of two parts: the first part measures whether each image belongs confidently to a single class of an ImageNet-1K pre-trained image classifier [98] and the second part measures how well the generated images capture diverse classes.
Precision and Recall [99]. The real and generated images are first converted to non-parametric representations of the manifolds using k-nearest neighbors, on which the Precision and Recall can be computed. Precision is the probability that a randomly generated image from estimated generated data manifolds falls within the support of the manifolds of estimated real data distribution. Recall is the probability that a random real image falls within the support of generated data manifolds. Thus, precision measures the general quality and fidelity of the generated images, and recall measures the coverage and diversity of the generated images.
@@ C.3 Control-to-Image Generation
Figure C: Performance of conditional generation for different condition types. We first introduce two baseline methods - ControlNet [1] and T2I-Adapter [2] to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser [96] implementation (for a fair comparison, ImageNet-pretrained LDM [6] is used as the base model). The performance is evaluated on the ImageNet validation set with our created pseudo labels.
Figure D: Performance of conditional generation for different condition types. We first introduce two baseline methods - ControlNet [1] and T2I-Adapter [2] to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser [96] implementation (for a fair comparison, ImageNet-pretrained LDM [6] is used as the base model). The performance is evaluated on the ImageNet validation set with our created pseudo labels.

### v2

@@ 4.1 Evaluation Settings
We conduct all the experiments on the ImageNet Deng et al. (2009) dataset. To incorporate pixel-level controls, we leverage state-of-the-art image understanding models to pseudo-label the images. Specifically, we label entity masks Kirillov et al. (2023), canny Canny (1986), depth Ranftl et al. (2020) and normal Vasiljevic et al. (2019) for both training and validation sets. This takes 500 Tesla V100 for about 4 days. We will release the pseudo-labeled datasets to facilitate the community to further explore conditional image generation.
@@
We utilize Fréchet Inception Distance (FID) Heusel et al. (2017), Inception Score (IS) Salimans et al. (2016), Precision, and Recall as metrics for assessing the quality of image generation. However, for the image-to-control prediction where ground truth is unavailable, we rely on qualitative visualization to demonstrate the perceptual quality.
@@
We follow VAR Tian et al. (2024) to use a GPT-2 Radford et al. (2019) style transformer with adaptive normalization Zhang et al. (2018). A transformer layer depth from 12 to 30 is explored. We leverage the pre-trained VAR tokenizer Tian et al. (2024) to tokenize both image and control. We initialize the model with the weights from VAR Tian et al. (2024) to shorten the training process. For each depth, we train the model for 30 epochs with an Adam optimizer. We follow the same learning rate and weight decay as VAR. During training, we sample each control type uniformly. To apply the classifier-free guidance, we replace class and control type conditions with empty tokens with 0.1 probability. We train the model with $\textrm{batchsize}=128$ for all the experiments. During inference, we utilize top-$k$ top-$p$ sampling with $k=900$ and $p=0.96$. We utilize $256\times 256$ image size for all experiments. For simplicity, we leverage $\gamma_{cls}=\gamma_{typ}=\gamma_{pix}$ for all the experiments.
@@ 4.2 Performance Analysis
Furthermore, we compare the image FID with pure image generation model VAR Tian et al. (2024) in Tab 1. We notice that ControlVAR shows a slight performance degradation compared to VAR which can be due to the difficulty enrolled to incorporate additional controls. As the model size increases, we notice the performance gap shrinks, indicating joint modeling of image and control may require more network capacity compared to image-only modeling.
Figure 7: Visualization of conditional image inpainting (given pixel-level control and partial image).
@@
We introduce two baseline methods - ControlNet Zhang et al. (2023) and T2I-Adapter Mou et al. (2023) to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser von Platen et al. (2022) implementation (for a fair comparison, ImageNet-pretrained LDM Rombach et al. (2021) is used as the base model). We compare the image generation quality in terms of FID and speed in Fig. 8. We notice that ControlVAR achieves obvious superior FID compared to baselines. We evaluate the inference speed with $\mathrm{batchsize}=1$ on a single H100 GPU. We notice that ControlVAR inference is at least 5 times faster than the compared methods. We further explore the generation capability with different model sizes as shown in Fig. 8 (b), we notice that the model’s generation capability keeps improving as the model size increases. Similar to the joint image-control generation, we notice that canny-conditioned generation shows an inferior performance due to its complex pattern.
@@
We further compare ControlVAR with conditional AR models - VQ-GAN Esser et al. (2021) and IQ-VAE Zhan et al. (2022). We finetune ControlVAR on ADE 20K for 1 epoch and report the FID of the generated images in Tab 2. ControlVAR demonstrates superior performance compared to previous AR methods.
@@ 5 Conclusion
In spite of ControlVAR’s high performance on image generation with heterogeneous pixel-level controls, it does not support text prompts and therefore cannot be directly leveraged with natural language guidance. Developing text-guided capability can be achieved by replacing the class token with the language token, e.g., CLIP token Radford et al. (2021), which is left as our future focus.
@@ Appendix A Dataset Creation
We conduct all the experiments on the ImageNet Deng et al. (2009) dataset. To incorporate pixel-level controls, we leverage state-of-the-art image understanding models to pseudo-label the images. Specifically, we label entity masks Kirillov et al. (2023), canny Canny (1986), depth Ranftl et al. (2020) and normal Vasiljevic et al. (2019) for both training and validation sets. This takes 500 Tesla V100 for about 4 days. We demonstrate the label number after filtering in Tab A. In addition, we also manually check the quality of the pseudo labels. We show a visualization of the generated datasets in Fig. A. We notice that image understanding models predict reasonable results on ImageNet images.
@@ Appendix B Discussion of the Teacher Forcing Guidance
Inspired by the classifier-free guidance Ho & Salimans (2022) from diffusion models, we empirically find a similar form of guidance that can be used for autoregressive sample conditional images based on teacher forcing. In this section, we aim to analyze the spirit of classifier-free guidance (CFG) and analogy it to our teacher-forcing guidance (TFG).
@@ B.1 Classifier-free Guidance
where $\gamma_{C},\gamma_{c},\gamma_{c_{t}}$ are the guidance scales that are used to adjust the amplitude to apply the conditional guidance. $\emptyset$ denotes leveraging a special empty token to replace the original token to disable the additional conditional information Ho & Salimans (2022).
@@ B.2 Teacher Forcing Guidance
Classifier-free guidance has been proven to be effective in AR models Chang et al. (2023); Tian et al. (2024) which take the same form as diffusion models as
@@
where $p(I,C|c,c_{t})$ and $p(I,C|c,c_{t})$ can be found from the output of ControlVAR. We follow previous works Tian et al. (2024); Esser et al. (2021) to ignore the constant probabilities $p(c)$. By rewriting all terms with Baysian rule, we have
@@ C.1 Details of Evaluation Metrics
Fréchet Inception Distance (FID) Heusel et al. (2017). FID measures the distance between real and generated images in the feature space of an ImageNet-1K pre-trained classifier Szegedy et al. (2016), indicating the similarity and fidelity of the generated images to real images.
Inception Score (IS) Salimans et al. (2016). IS also measures the fidelity and diversity of generated images. It consists of two parts: the first part measures whether each image belongs confidently to a single class of an ImageNet-1K pre-trained image classifier Szegedy et al. (2016) and the second part measures how well the generated images capture diverse classes.
Precision and Recall Kynkäänniemi et al. (2019). The real and generated images are first converted to non-parametric representations of the manifolds using k-nearest neighbors, on which the Precision and Recall can be computed. Precision is the probability that a randomly generated image from estimated generated data manifolds falls within the support of the manifolds of estimated real data distribution. Recall is the probability that a random real image falls within the support of generated data manifolds. Thus, precision measures the general quality and fidelity of the generated images, and recall measures the coverage and diversity of the generated images.
@@ C.3 Control-to-Image Generation
Figure C: Performance of conditional generation for different condition types. We first introduce two baseline methods - ControlNet Zhang et al. (2023) and T2I-Adapter Mou et al. (2023) to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser von Platen et al. (2022) implementation (for a fair comparison, ImageNet-pretrained LDM Rombach et al. (2021) is used as the base model). The performance is evaluated on the ImageNet validation set with our created pseudo labels.
Figure D: Performance of conditional generation for different condition types. We first introduce two baseline methods - ControlNet Zhang et al. (2023) and T2I-Adapter Mou et al. (2023) to compare the conditional generation capability. We train both baselines on the same datasets as ours with the Diffuser von Platen et al. (2022) implementation (for a fair comparison, ImageNet-pretrained LDM Rombach et al. (2021) is used as the base model). The performance is evaluated on the ImageNet validation set with our created pseudo labels.